james has 4and3/4 feet of rope he plans to cut iff1and1/2 feet of rope how much will be left
Answer by rm29924(97)

4 3/4 - 1 1/2
change fractions to improper fractions
19/4 - 3/2
change it so they have the same denominator
19/4 - 6/4 = 13/4
change the answer back into a mixed fraction
3 1/4

check answer:
1 1/2 + 3 1/4=
3/2 + 13/4=
6/4 + 13/4=19/4
4 3/4